OpenStudy (anonymous):

A circle is drawn through the vertices of a square with sides of 5 inches. If a dart is thrown at random into the circle, what is the probability it will NOT lie in the square. Use pythagoreean theorem to find the radius of the circle.

OpenStudy (anonymous):

probability = 1 - 25/pi*50 = 1 - 1/2pi = (2 pi -1 ) /2pi

OpenStudy (anonymous):

can u draw it out

OpenStudy (anonymous):

diameter of cirlcle = sqrt (5^2 + 5^2) = sqrt50 area of the circle = pi * (sqrt50/2)^2 sq ins area of square = 25 sq ins area not in square = pi * (sqrt50/2)^2 - 25 probability that dart is not in square = [pi * (sqrt50/2)^2 - 25] / pi * (sqrt50/2)^2

OpenStudy (anonymous):

jimmyrep did it for me :P

OpenStudy (anonymous):

soo whats the answer

OpenStudy (anonymous):

i told you the answer already Ans = 1 - (1/2pi )